# Improving planning and MBRL with temporally-extended actions

### v1

@@ Appendix A Iterative Primitive Dynamics Function
2:$\text{repeats}=\lfloor{\nicefrac{{\delta t_{k}}}{{\delta_{t}^{\text{env}}}}}\rfloor$
@@
4:     $s_{k}=f(s_{k},a_{k},\delta_{t}^{\text{env}})$triangleright Control $n$
@@
6:$s_{k+1}=f(s_{k},a_{k},\delta t_{k}~{}\texttt{mod}~{}\delta_{t}^{\text{env}})$triangleright Control $\delta_{t}$
@@
Appendix B Connection to Ni and Jang [2022]
Ni and Jang [2022] propose to learn a policy that outputs the action variables as well as the time scale. They have a similar looking objective. In this section, we compare our objective to theirs. First, we repeat Equation 1 which computes the returns of a trajectory $\tau$ using the proposed framework:
$J(\tau)=\sum_{k=1}^{L(\tau)}\gamma^{e_{<k}}\sum_{t=1}^{e_{k}}\gamma^{t-1}\mathcal{R}_{(e_{<k}+t)}$
A similar objective is used by Ni and Jang [2022]. However, it is important to note that they control the timescale while we keep the timescale fixed and control the action duration. Let $\delta_{k}\in[\delta_{\min},\delta_{\max}]$ be the timescale associated with the action at decision point $k$. For simplicity, they assume that the timescale is in integers (in physical unit of seconds), which means that $\delta_{k}$ is also the number of execution steps associated with the action. Using this, their objective reduces to
where $\mathcal{R}(s_{e_{<k}+1},a_{k})\delta_{k}$ is a linear approximation of the reward function. So, the objective used by Ni and Jang [2022] can be viewed as an approximation of Equation 1 where the reward function due to a temporally-extended action has been replaced by a linear approximation.
Appendix C Environment Overview
In this section, we provide an overview of the environments used in our experiments. Table A1 lists the dimensionality of the observation and the action space of the environments, the reward functions along with the episode length (which is equivalent to the maximum number of primitive steps allowed).
@@ Appendix C Environment Overview
[…]
@@
Hopper 11 3 1000 $(\text{sigmoid}(\dot{x_{t}})\text{ if }\dot{x_{t}}\leq 2\text{ else }1)\times\text{healthy reward}$
@@ C.1 MuJoCo environments
Reward functions : The default reward functions in MuJoCo are not bounded. For the MAB framework, we require a bounded reward function. The problem is amplifed in Ant, Half Cheetah, Hopper and Walker, where the rewards over an episode can be very large. So, for these environments, we adapt the bounded reward functions from DeepMind Control Suite [Tunyasuvunakool et al., 2020], so that the agents always get a reward between 0 and 1, with the specifics depending on the environment.
Maximum episode lengths : MuJoCo environments have a predefined value of frame-skip. The default episode length is with respect to this frame-skip. For example, Half Cheetah has a frame-skip of 5 and a episode length of 1000. An equivalent episode in Half Cheetah with a frame-skip of 1 should be of length 5000. However, we limit the episode to 1000 as otherwise $A_{\text{STD}}$ would have taken too long to train. Even with episode lengths of 1000, $A_{\text{STD}}$ requires more than 40 hours to train (Table A3).
Appendix D Experimental Details
Our framework has two main hyperparameters - range of action duration, which is specified by $\delta t_{\min}$ and $\delta t_{\max}$, and planning horizon. Rather than controlling both these values, $\delta t_{\min}$ is always set to $\delta_{t}^{\text{env}}$, and modifying $\delta t_{\max}$ controls the range. For the environments Cartpole and Mountain Car, we use values of planning horizon from prior work for the standard framework, and adjust the depth for $A_{\text{TE}}$ by exploring related values. We cannot do this for MuJoCo-based environments since we set the frameskip parameter to 1. Rather we perform a search over some potential values for the planning horizon for the standard framework and $\delta t_{\max}$ and planning horizon for our framework and choose the configuration with the best performance. The other hyperparameters related to online training have been borrowed from Chua et al. [2018].
@@ Appendix D Experimental Details
Details for Dubins Car environment : The default timescale for the environment is $0.2$ while $\delta t_{\max}$ for our formulation is 20. Setting $\delta t_{\max}$ to such a large value allows us to have the same value across maps and just allow more time for optimization. The planning horizon needs to be tuned for different maps. Table A2 contains the planning horizon for the maps used in the experiments. Note that for our experiment with the cave-mini map, we reduce $\delta t_{\max}$ to prevent the agent from completing the map in a few decision steps.
@@
Algorithm A2 Planning using CEM with ensemble of dynamics model.
1:for time $t=0$ to $H$ do
2:     sample $n$ action sequences $\{a^{i}_{t},a^{i}_{t+1},\dots,a^{i}_{t+D}\}$
3:     for each sampled action sequence do
4:         obtain state sequence by propagating particles using $s^{i}_{t+1}=\tilde{f}(s^{i}_{t},a^{i}_{t})$
5:         evaluate the action sequence using $\sum_{t=1}^{D}\sum_{p=1}^{P}\nicefrac{{r(s^{p}_{t},a^{p}_{t})}}{{P}}$
6:         choose the top $k$ action sequences to update CEM()($\cdot$)( ⋅ ) distribution.
7:     end for
8:     Execute first action $a$ from optimal action sequence.
9:end for
Model learning : For MBRL, we use the same model architecture as Chua et al. [2018]. We learn an ensemble of 5 models, where each model is a fully connected neural network. For Ant, Half Cheetah and Hopper, the model has 4 hidden layers while for all other environments, it has 3 hidden layers. Each hidden layer has 200 neurons. The learning rates for each environment are given in Table A2.
Appendix E Computational Resources
Each MBRL experiment was performed on a single node with a single GPU (using a mix of V100 and A100), single CPU (AMD EPYC 7742) and 64GB of RAM. The time required for training for different algorithms are in Table A3. We note that although the experiments were done on two different kinds of GPUs, the running times for both of them are comparable.
@@ Appendix E Computational Resources
In Ant, Half Cheetah, Hopper and Walker, $A_{\text{TE}}$ is significantly faster than $A_{\text{STD}}$. The dynamic variant of our proposed algorithm takes slightly longer than the fixed variant as the rollout length of the dynamic variant is longer on average than the rollout length of the fixed variant.
Appendix F Additional Planning Experiments
F.1 Experiments with IPC Mountain Car
@@
We vary the planning horizon of the agents and compare the performances across instances in Table A4. For the easier problems, a small planning horizon is sufficient, but for the difficult instances, a deeper search is required. The performance of the planner remains relatively stable as the planning horizon increases.
@@ D Instance 1 Instance 2 Instance 3 Instance 4 Instance 5
F.2 Experiments with U-shaped Maps in Dubins Car
@@
To be successful in this setting, we observe that $A_{\text{STD}}$ requires a planning horizon of 1000 and $A_{\text{TE}}$ requires a planning horizon of 75. Both the agents use 10,000 samples and 50 optimization steps for each decision. $A_{\text{TE}}$ succeeds but not in all runs. We look at the failure cases for $A_{\text{TE}}$ and have an interesting observation. For planning horizon of 75 and 100, the failure is due to the fact that the agent runs out of primitive actions, while for planning horizon 50, the failing scenario corresponds to the agent not being able to find a path out of the obstacles region. This means that if we allow the episodes to run for longer, the former failure can be mitigated but the latter cannot. The detailed results are in Table A5 and the failure cases for $A_{\text{TE}}$ are shown in Figure A2.
@@
Next, we make the problem difficult by augmenting the action space with 100 dummy action variables. As the agents are unaware that these variables don’t contribute to the reward or the dynamics, they still need to search over these variables. We setup the experiment similar to the previous one by keeping the number of samples to 10000 and varying the planning horizon. For $A_{\text{STD}}$, almost all the runs fail - either because it does not find a path around the obstacle, or crashes due to huge memory requirements. In order to reduce the memory requirements, we perform a second experiment, fixing the planning horizon to 1000 and varying the number of samples. Reducing the number of samples fixes the issue of memory requirements, but does not help the agent identify a path. In contrast, the performance of $A_{\text{TE}}$ is mostly unaffected by this addition of dummy action variables. The detailed results are in Table A6. $A_{\text{TE}}$ uses the same configuration as the earlier experiment and achieves similar performance. As observed earlier, the failure case of $A_{\text{TE}}$ for planning horizon of 50 is due to the fact that it cannot find a path around the obstacles, while the failure case for planning horizon of 100 is because it runs out of primitive steps.
Appendix G Additional MBRL results
How do the action repeats due to $A_{\text{TE}}$ vary across training iterations?
@@ Appendix G Additional MBRL results
While the histogram in Figure 4 gives us a good sense of the distribution of discretized action duration $(\nicefrac{{\delta t}}{{\delta_{t}^{\text{env}}}})$ chosen by $A_{\text{TE}}$ in one episode after the training has been completed, it does not capture the trend across time. As the average action repeat for $A_{\text{TE}}(\text{D})$ can have a large range, in Figure A3 , we plot the log of average action repeat instead. $A_{\text{STD}}$ has an action repeat of 1 as it works with primitive actions. For $A_{\text{TE}}(\text{F})$, the average action repeat is mostly stable. In Cartpole, Reacher and Pusher, the agent has larger action repeats (or takes large values of $\delta t$) in the beginning before realizing that smaller action repeats work better. The more interesting plots are that of $A_{\text{TE}}(\text{D})$ as it has to figure out the correct range of $\delta t$. From the plots, we see that the average action repeats decreases gradually before stabilizing around the average action repeat used by $A_{\text{TE}}(\text{F})$. Note that larger values of average action repeats does not necessarily mean smaller number of decision points. $A_{\text{TE}}(\text{D})$ has to choose $\delta t_{\max}$ candidates from a list of exponentially spaced candidates and a large $\delta t_{\max}$ selection can easily skew the average action repeat.
Appendix H Ablations
H.1 Single model vs. Separate models
@@ Appendix H Ablations
As shown in Figure A4, we find that using the UCB heuristic, the agent performs better if it learns a separate dynamics model for each bandit.

### v2

@@ Appendix B Connection to Ni and Jang [2022]
[…]
@@
4:  $s_{k}=f(s_{k},a_{k},\delta_{t}^{\text{env}})$ $\triangleright$ Control $n$
@@
[…]
@@
Appendix C Iterative Primitive Dynamics Function
In most cases we have access to the primitive dynamics function $(f)$. A trivial way of making $f$ work with temporally-extended actions  is to wrap it in a loop. For the computation to be exact, $f$ should be explicitly dependent on time. If it is implicit, performing the computation at line 5 of Algorithm˜A1 will not be possible and the algorithm computes the number of action-repeats instead.
@@ Appendix C Iterative Primitive Dynamics Function
Appendix D Environment Overview
In this section, we provide an overview of the environments used in our experiments. Table˜A1 lists the dimensionality of the observation and the action space of the environments, the reward functions along with the episode length (which is equivalent to the maximum number of primitive steps allowed).
@@ Appendix D Environment Overview
Pusher 23 7 150 $0.5\times\text{dist(finger, obj)}-\text{dist(object, goal)}-0.1a_{t}^{2}$
@@
[…]
@@ D.1 MuJoCo environments
Reward functions : The default reward functions in MuJoCo are not bounded. For the MAB framework, we require a bounded reward function. The problem is amplifed in Ant, Half Cheetah, Hopper and Walker, where the rewards over an episode can be very large. Therefore, for these environments, we adapt the bounded reward functions from DeepMind Control Suite [Tunyasuvunakool et al., 2020], so that the agents always get a reward between 0 and 1, with the specifics depending on the environment.
Maximum episode lengths : MuJoCo environments have a predefined value of frame-skip. The default episode length is with respect to this frame-skip. For example, Half Cheetah has a frame-skip of 5 and a episode length of 1000. An equivalent episode in Half Cheetah with a frame-skip of 1 should be of length 5000. However, we limit the episode to 1000 as otherwise $A_{\text{STD}}$ would have taken too long to train. Even with episode lengths of 1000, $A_{\text{STD}}$ requires more than 40 hours to train (Table˜A3).
Appendix E Experimental Details
Among the set of hyperparameters in our framework, the two crucial ones are the range of action duration, which is specified by $\delta t_{\min}$ and $\delta t_{\max}$, and the planning horizon. Rather than controlling both these values, $\delta t_{\min}$ is always set to $\delta_{t}^{\text{env}}$, and modifying $\delta t_{\max}$ controls the range.
Details for Cartpole, Mountain Car and MuJoCo : For the environments Cartpole and Mountain Car, we use values of planning horizon from prior work for the standard framework, and adjust the depth for $A_{\text{TE}}$ by exploring related values. We could not do this for MuJoCo-based environments since we set the frameskip parameter to 1. Rather we searched over some potential values for the planning horizon for the standard framework, and $\delta t_{\max}$ and planning horizon for our framework and chose the configuration with the best performance. The other hyperparameters related to online training have been borrowed from Chua et al. [2018].
@@ Appendix E Experimental Details
Details for Dubins Car environment : The default timescale for the environment is $0.2$ while $\delta t_{\max}$ for our formulation is 20. Setting $\delta t_{\max}$ to such a large value allows us to have the same value across maps and just allow more time for optimization. The planning horizon needs to be tuned for different maps. Table˜A2 contains the planning horizon for the maps used in the experiments. Note that for our experiment with the cave-mini map, we reduce $\delta t_{\max}$ to prevent the agent from completing the map in a few decision steps.
Model learning : For MBRL, we use the same model architecture as Chua et al. [2018]. We learn an ensemble of 5 models, where each model is a fully connected neural network. For Ant, Half Cheetah and Hopper, the model has 4 hidden layers while for all other environments, it has 3 hidden layers. Each hidden layer has 200 neurons. The learning rates for each environment are given in Table˜A2.
@@
[…]
2:for $i=1$ to number of updates do
3:  Sample NN action sequences $\{a^{N}_{1},a^{N}_{2},\dots,a^{N}_{D}\}$ using $\mu_{a},\sigma_{a}$.
4:  for $n=1$ to NN do
5:   Initialize $P$ particles = $\{a^{n,P}_{1},a^{n,P}_{2},\dots,a^{n,P}_{D}\}$.
6:   for $p=1$ to $P$ do
7:     //propagate each particle using one model from ensemble
8:     $j=p\texttt{ mod }K$
9:     Propagate particle $p$ using $s^{n,p}_{t+1}=\hat{f}_{j}(s^{n,p}_{t},a^{n,p}_{t})$ and compute aggregate reward.
10:   end for
11:   Compute mean reward for action sequence using $\sum_{t=1}^{D}\sum_{p=1}^{P}\nicefrac{{r(s^{n,p}_{t},a^{n,p}_{t})}}{{P}}$
12:  end for
13:  Choose the top $k$ action sequences and recompute $\mu_{a},\text{var}_{a}$
14:end for
15:Execute first action from the optimal action sequence.
[…]
$\mu_{a}$ $=\frac{1}{k}\sum_{a\in\mathcal{K}}a$ (5)
$\text{var}_{a}$ $=\frac{1}{k}\sum_{a\in\mathcal{K}}(a-\mu_{a})^{2}$ (6)
[…]
Appendix F Computational Resources
Each MBRL experiment was performed on a single node with a single GPU (using a mix of V100 and A100), single CPU (AMD EPYC 7742) and 64GB of RAM. The time required for training for different algorithms are in Table˜A3. We note that although the experiments were done on two different types of GPUs, the running times for both of them are comparable.
@@ Appendix F Computational Resources
In Ant, Half Cheetah, Hopper and Walker $A_{\text{TE}}$ is significantly faster than $A_{\text{STD}}$. The dynamic variant of our proposed algorithm takes slightly longer than the fixed variant as the rollout length of the dynamic variant is longer on average than the rollout length of the fixed variant.
Appendix G Additional Planning Experiments
G.1 Experiments with IPC Mountain Car
@@
We vary the planning horizon of the agents and compare the performances across instances in Table˜A4. For the easier problems, a small planning horizon is sufficient, but for the difficult instances, a deeper search is required. The performance of the planner remains relatively stable as the planning horizon increases.
@@ D Instance 1 Instance 2 Instance 3 Instance 4 Instance 5
G.2 Experiments with U-shaped Maps in Dubins Car
@@
To be successful in this setting, we observe that $A_{\text{STD}}$ requires a planning horizon of 1000 and $A_{\text{TE}}$ requires a planning horizon of 75. Both the agents use 10,000 samples and 50 optimization steps for each decision. $A_{\text{TE}}$ succeeds but not in all runs. We look at the failure cases for $A_{\text{TE}}$ and have an interesting observation. For planning horizon of 75 and 100, the failure is due to the fact that the agent runs out of primitive actions, while for planning horizon 50, the failing scenario corresponds to the agent not being able to find a path out of the obstacles region. This means that if we allow the episodes to run for longer, the former failure can be mitigated but the latter cannot. The detailed results are in Table˜A5 and the failure cases for $A_{\text{TE}}$ are shown in Figure˜A2.
@@
Next, we make the problem difficult by augmenting the action space with 100 dummy action variables. As the agents are unaware that these variables don’t contribute to the reward or the dynamics, they still need to search over these variables. We set up the experiment similar to the previous one by keeping the number of samples to 10000 and varying the planning horizon. For $A_{\text{STD}}$, almost all the runs fail - either because it does not find a path around the obstacle, or crashes due to huge memory requirements. In order to reduce the memory requirements, we perform a second experiment, fixing the planning horizon to 1000 and varying the number of samples. Reducing the number of samples fixes the issue of memory requirements, but does not help the agent identify a path. In contrast, the performance of $A_{\text{TE}}$ is mostly unaffected by this addition of dummy action variables. The detailed results are in Table˜A6. $A_{\text{TE}}$ uses the same configuration as the earlier experiment and achieves similar performance. As observed earlier, the failure case of $A_{\text{TE}}$ for planning horizon of 50 is due to the fact that it cannot find a path around the obstacles, while the failure case for planning horizon of 100 is because it runs out of primitive steps.
Appendix H Additional MBRL results
H.1 How do the action repeats due to $A_{\text{TE}}$ vary across training iterations?
@@
While the histogram in Figure˜4 gives us a good sense of the distribution of discretized action duration $(\nicefrac{{\delta t}}{{\delta_{t}^{\text{env}}}})$ chosen by $A_{\text{TE}}$ in one episode after the training has been completed, it does not capture the trend across time. As the average action repeat for $A_{\text{TE}}(\text{D})$ can have a large range, in Figure˜A3 , we plot the log of average action repeat instead. $A_{\text{STD}}$ has an action repeat of 1 as it works with primitive actions. For $A_{\text{TE}}(\text{F})$, the average action repeat is mostly stable. In Cartpole, Reacher and Pusher, the agent has larger action repeats (or takes large values of $\delta t$) in the beginning before realizing that smaller action repeats work better. The more interesting plots are those of $A_{\text{TE}}(\text{D})$ as it has to identify the correct range of $\delta t$. From the plots, we see that the average action repeats decreases gradually before stabilizing around the average action repeat used by $A_{\text{TE}}(\text{F})$. Note that larger values of average action repeats does not necessarily mean smaller number of decision points. $A_{\text{TE}}(\text{D})$ has to choose $\delta t_{\max}$ candidates from a list of exponentially spaced candidates and a large $\delta t_{\max}$ selection can easily skew the average action repeat.
Appendix I Ablations
I.1 Single model vs. Separate models
@@ Appendix I Ablations
As shown in Figure˜A4, we find that using the UCB heuristic, the agent performs better if it learns a separate dynamics model for each bandit arm.
I.2 Which is a better approximation for $F$ - $\hat{F}_{\text{TE}}$ or $\hat{F}_{\text{IP}}$?
While we proposed to use $\hat{F}_{\text{TE}}$ to approximate the temporally-extended dynamics function $F$, one could also consider using $\hat{F}_{\text{IP}}$ which approximates the one-step dynamics function $f$ and then iterates over it. While we know that $\hat{F}_{\text{TE}}$ leads to better time complexity than $\hat{F}_{\text{IP}}$, it is interesting to explore how it compares in terms of prediction quality.
To answer this question, we perform two experiments with random action sequences and compare the prediction quality of $\hat{F}_{\text{TE}}$ and $\hat{F}_{\text{IP}}$. In both the experiments, we sample random action sequences of length 500 and compute the mean-squared error (MSE) between the actual states observed upon taking the action sequences in the environment and the states as predicted by the learned model. The difference between the two experiments lies in the type of action sequences being sampled. For the first experiment, we randomly sample one-step action sequences which implicitly correspond to duration $\delta t_{\min}$, while for the second experiment, we randomly sample temporally-extended action sequences with action duration sampled from $\mathcal{U}(\delta t_{\min},\delta t_{\max})$. Each experiment was repeated 10 times in order to get a better estimate and the mean results are shown in Table˜A7.
Random one step sequences Random temporally extended sequences
Env $\delta t_{\min}$ $\delta t_{\max}$ $\hat{F}_{\text{TE}}$ $\hat{F}_{\text{IP}}$ $\hat{F}_{\text{TE}}$ $\hat{F}_{\text{IP}}$ $\hat{F}_{\text{IP}}$ without outliers
Cartpole 1 10 0.0107 $2.517\times 10^{-7}$ 0.0184 0.065 0.065
Ant 0.01 0.07 0.126 0.0149 0.298 $4.525\times 10^{10}$ 0.60
Half Cheetah 0.01 0.07 0.30 0.04 1.752 2.526 2.526
Table A7: Comparison between the prediction error (in terms of MSE) of $\hat{F}_{\text{TE}}$ and $\hat{F}_{\text{IP}}$ on random action sequences.
In Ant, when predicting using temporally-extended action sequences, in 4 out of 10 experiments, the MSE due to $\hat{F}_{\text{IP}}$ was extremely high. We hypothesize that the states observed during these experiments were probably outside the range of examples seen during the training of the one-step model that is used by $\hat{F}_{\text{IP}}$. $\hat{F}_{\text{IP}}$ without outliers reports the MSE after removing these 4 experiments.
As can be observed, for one-step sequences $\hat{F}_{\text{IP}}$ has a lower MSE but for temporally-extended sequences, $\hat{F}_{\text{TE}}$ is clearly the better choice. This is not surprising since using temporally-extended action sequences with $\hat{F}_{\text{IP}}$ requires iterating over the one-step model multiple times which can lead to compounding errors. Thus, $\hat{F}_{\text{TE}}$ is a better choice both in terms of time complexity and prediction quality when using temporally-extended action sequences.